MICROWAVE SPECTRA OF FLUORINATED PROPIONIC ACIDS AND THEIR HYDRATES

DANIEL A. OBENCHAIN, G. S. GRUBBS II, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; STEPHEN A. COOKE, School of Natural and Social Sciences, Purchase College, SUNY, 735 Anderson Hill Road, Purchase, NY 10577; AGAPITO SER-RATO III and \underline{WEI LIN}, Department of Chemistry and Environmental Science, University of Texas at Brownsville, Brownsville, TX 78520.

The microwave spectra of three fluorinated propionic acids, namely pentafluoropropionic acid, 2,2,3,3-tetrafluoropropionic acid, and 2,3,3,3-tetrafluoropropionic acid have been observed and assigned. The spectra of pentafluoropropionic acid and its monohydrate were reported in this meeting last year. The conformational analysis of 2,2,3,3-tetrafluoropropionic acid and 2,3,3,3-tetrafluoropropionic acid will be compared to that of pentafluoropropionic acid. The structures of the three monohydrates will be discussed in terms of hydrogen bonding.